OpenStudy (anonymous):

[10.02] Solve the proportion 2 times x all over 9 times x minus 8 equals 2 over x.. What is the value of x?

OpenStudy (anonymous):

Is the equation like this : \[\Large\frac{2x}{9x-8}=\frac{2}{x}\] ??

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

nd there's two values for x

OpenStudy (anonymous):

To solve this equation we use the cross : \[\Large \frac{\color{red}a}{\color{blue}b}=\frac{\color{blue}c}{\color{red}d}\Longrightarrow \color{red}{ad}=\color{blue}{bc}\]

OpenStudy (anonymous):

Yeah I knew tht

OpenStudy (anonymous):

Good ! So you will get : \[\Large 2x^2=18x-16\] So, by dividing over 2 we get : \[\Large x^2-9x+8=0\] This can be factored like this : \[\Large (x-8)(x-1)=0\] So : \[\Large x-8=0\quad \text{ or } \quad x-1=0\] So : \[\Large x=8\quad \text{ or } \quad x=1\]

OpenStudy (anonymous):

I get uber confused with factoring. lol thanks c:

OpenStudy (anonymous):

To factor the expression : \[\Large x^2-9x+8\] we have to look for two numbers a and b such that : \[\Large a+b=-9 \text{ and } ab=8\]

OpenStudy (anonymous):

Yeah I think that is what throws me off

OpenStudy (anonymous):

I have 2 problems I am stuck on would you mind helping me out I really need to finish this test so I can take my exam.